i am trying to determine if my son and I solved the following equation correctly:
Questions: Find the slope of the line passing through (-5/4,2) and (1/4,7/2)

Answer by Nate(3500)

Slope = (y2 - y1)/(x2 - x1) where your points are defined as: (x1,y1) and (x1,y1)
Slope = ((7/2) - 2)/((1/4) + (5/4))
Slope = (3.5 - 2)/(0.25 + 1.25)
Slope = (1.5)/(1.5) = 1/1